# Persuade Me if You Can: A Framework for Evaluating Persuasion Effectiveness and Susceptibility Among Large Language Models

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate persuasive capabilities that rival human-level persuasion. While these capabilities can be used for social good, they also present risks of potential misuse. Moreover, LLMs’ susceptibility to persuasion raises concerns about alignment with ethical principles. To study these dynamics, we introduce Persuade Me If You Can (PMIYC), an automated framework for evaluating persuasion through multi-agent interactions. Here, Persuader agents engage in multi-turn conversations with the Persuadee agents, allowing us to measure LLMs’ persuasive effectiveness and their susceptibility to persuasion. We conduct comprehensive evaluations across diverse LLMs, ensuring each model is assessed against others in both subjective and misinformation contexts. We validate the efficacy of our framework through human evaluations and show alignment with prior work. PMIYC offers a scalable alternative to human annotation for studying persuasion in LLMs. Through PMIYC, we find that Llama-3.3-70B and GPT-4o exhibit similar persuasive effectiveness, outperforming Claude 3 Haiku by 30%. However, GPT-4o demonstrates over 50% greater resistance to persuasion for misinformation compared to Llama-3.3-70B. These findings provide empirical insights into the persuasive dynamics of LLMs and contribute to the development of safer AI systems. [^1] https://beyzabozdag.github.io/PMIYC
Persuade Me if You Can: A Framework for Evaluating Persuasion Effectiveness and Susceptibility Among Large Language Models
Nimet Beyza Bozdag, Shuhaib Mehri, Gokhan Tur, Dilek Hakkani-Tür University of Illinois Urbana-Champaign {nbozdag2, mehri2, gokhan, dilek}@illinois.edu
@@ 1 Introduction
Figure 1: A $t$-turn interaction between a Persuader and a Persuadee in PMIYC. The Persuader can addresses the Persuadee’s concerns, while the Persuadee reports its agreement scores.
Large Language Models (LLMs) have demonstrated remarkable capabilities across various domains and are now integral to numerous real-world applications (OpenAI, 2024c; Anthropic, 2024). Among these capabilities, persuasion stands out as particularly noteworthy, with state-of-the-art LLMs exhibiting persuasive skills comparable to humans (Durmus et al., 2024; OpenAI, 2024c). Persuasive LLMs have been applied for social good, promoting public health and prosocial behaviors while adhering to ethical, non-coercive principles (Wang et al., 2019; Furumai et al., 2024; Karinshak et al., 2023; Kampik et al., 2018). However, as LLMs become increasingly persuasive, they also pose risks. These models can be exploited to manipulate individuals, spread misinformation, and influence public opinion in harmful ways (Singh et al., 2024; Salvi et al., 2024; Simchon et al., 2024; Hackenburg and Margetts, 2024).
Beyond influencing others, LLMs themselves are susceptible to persuasion, making them vulnerable to manipulation that may lead to violations of alignment principles, generating toxic or misleading content, and reinforcing biases (Zeng et al., 2024; Xu et al., 2024). This presents a challenge in having LLMs resist harmful or inappropriate influence while remaining open to justified reasons to change their views (Stengel-Eskin et al., 2024).
@@
In this work, we introduce Persuade Me If You Can (PMIYC), an automated framework for evaluating persuasion in LLMs across diverse conversational setups. Building on recent studies (Xu et al., 2024; Stengel-Eskin et al., 2024), our framework simulates multi-agent interactions where a Persuader attempts to convince a Persuadee to agree with a claim. PMIYC enables the Persuader and Persuadee to engage in multi-turn conversations, exchanging arguments, addressing counterpoints, and influencing each other’s views. Throughout these interactions, the Persuadee continuously assesses the Persuader’s arguments and self-reports its level of agreement, allowing us to track changes in opinion and measure both persuasive effectiveness (how successfully an LLM persuades) and susceptibility to persuasion (how easily an LLM changes its stance).
Our round-robin evaluation, where each model takes turns acting as both the Persuader and the Persuadee against every other model, enables direct comparisons. Our evaluation spans a diverse set of models, including open-weight models (e.g., Llama-3.3-70B-Instruct) and closed-source models (e.g., GPT-4o, Claude 3 Opus). We assess single-turn persuasion, where the Persuader has only one chance to convince the Persuadee, as well as multi-turn persuasion, which allows for back-and-forth exchanges and deeper argumentation. Additionally, we examine persuasion across different contexts, including subjective claims (eg. "Cultured/lab-grown meats should be allowed to be sold") and misinformation claims (eg. "Canada is a part of the UK"), to evaluate how LLMs respond under varying conditions.
Our results reveal that persuasive effectiveness remains largely consistent across settings, with Llama-3.3-70B-Instruct emerging as the most effective persuader, closely followed by GPT-4o. However, susceptibility to persuasion varies based on the context and the number of conversational turns. Notably, GPT-4o and GPT-4o-mini exhibit up to 50% stronger resistance to persuasion in misinformation contexts but show greater flexibility with subjective claims. Persuasive effectiveness is greater in the first two attempts, and a Persuader becomes more effective when it more strongly agrees with the claim it advocates. By modeling persuasion as an interactive and iterative process, PMIYC provides a more realistic setting for studying persuasion as it occurs in conversations. The reliability of our framework is validated through human evaluations and comparative analyses with existing studies. Our findings align closely with human judgments of persuasiveness, reinforcing the validity of PMIYC as a scalable alternative to human assessment. Our key contributions are:
  • PMIYC, a framework for automatically evaluating both the persuasive effectiveness and susceptibility to persuasion of LLMs.
@@
Through this research, we aim to deepen the understanding of persuasive capabilities and susceptibility of LLMs, and to emphasize the safety risks associated with persuasion in AI systems. By providing a scalable and automated evaluation framework, we aim to contribute to the development of safer, more transparent, and ethically responsible LLMs for real-world applications.
@@ 2 Related Work
Recent studies suggest that LLMs exhibit human-like persuasiveness (Durmus et al., 2024; OpenAI, 2024c; Huang and Wang, 2023). Durmus et al. (2024)’s and OpenAI (2024c)’s experiments show that their models are as persuasive as humans. While some studies leverage the persuasive abilities of LLMs for social good Karinshak et al. (2023); Costello et al. (2024), others warn of risks like microtargeting, political influence, and ethical concerns Salvi et al. (2024); Simchon et al. (2024); Hackenburg and Margetts (2024). These findings underscore the growing urgency of understanding and regulating persuasive AI systems.
Evaluating Persuasion.
@@
Prior research has explored various methods to measure the persuasiveness of LLMs. Durmus et al. (2024) uses human evaluators to rank their agreement with a subjective claim before and after reading the model-generated arguments. While this approach provides valuable insights into human preferences, it is not highly scalable. Similarly, OpenAI uses human annotators by having them rank arguments and perform pairwise comparisons OpenAI (2024c). They also explore automated evaluation, through their MakeMeSay and MakeMePay role-playing frameworks. These frameworks, while innovative, are highly context-dependent and may not generalize well to real-world persuasion scenarios. Singh et al. (2024) introduce PersuasionArena and the task of transsuasion, where LLMs rewrite low-engagement tweets into more persuasive versions. Although PersuasionArena provides an automated framework for measuring persuasion, its focus is less conversational and does not fully address the interactive and manipulative aspects of persuasion. Breum et al. (2023) develop a Convincer-Skeptic scenario, in which an LLM attempts to persuade another LLM in the domain of climate change. They show that agents can mimic some of the persuasive dynamics and opinion changes observed in human discourse. Other studies have attempted to score persuasive language in pairs of LLM-generated text using a regression model (Pauli et al., 2024). Yet, these efforts focus on short text excerpts, limiting their applicability to longer, multi-turn persuasion.
LLM Susceptibility to Persuasion.
@@
While LLMs generating persuasive content raises ethical concerns, their susceptibility to persuasion is equally important. Some jailbreaking studies show that LLMs can be persuaded to generate harmful content Zeng et al. (2024) or accept and reinforce misinformation Xu et al. (2024). Furthermore, recent studies have shown that multi-turn jailbreak attempts can be particularly effective in circumventing safeguards Xu et al. (2024); Li et al. (2024); Russinovich et al. (2024). Inspired by these findings, we extend our framework to include the misinformation domain.
@@
Overall, prior research has primarily focused on single-turn persuasive attacks on LLMs, human-in-the-loop evaluation methods, less conversational or highly constrained scenarios, with limited exploration of multi-turn agent-to-agent persuasion. In contrast, we explore how LLMs interact with each other in a highly conversational setting across different domains and multi-turn scenarios, encompassing both beneficial and harmful persuasion.
@@ 3 Methodology
PMIYC is our framework for evaluating both the persuasive effectiveness of LLMs and their susceptibility to persuasion. [^2] Code adapted from NegotiationArena Bianchi et al. (2024). Experiments consist of multi-agent interactions, where one agent acts as a Persuader (ER) and the other acts as a Persuadee (EE). The two engage in a structured exchange, where the Persuader agent is given the objective of convincing the Persuadee to agree with a given claim. The Persuadee engages with the Persuader and re-assesses its stance on the claim throughout the conversation.
@@ 3.1 Persuasive Conversation Generation
Conversation generation involves simulating a dialogue between two agents: a Persuader and a Persuadee. The objective is to track how the Persuadee’s stance on a given claim $C$ evolves over the course of a $t$-turn interaction. The conversation begins with the Persuadee expressing its initial agreement with the claim by providing both a message and an agreement score (see Turn 1 in Figure 1). The agreement score quantifies the model’s stance on claim $C$ using a five-point Likert scale: Completely Oppose (1), Oppose (2), Neutral (3), Support (4), and Completely Support (5). This scale enables a more granular analysis of opinion shifts beyond a binary agree/disagree classification.
After observing the Persuadee’s initial stance, the Persuader is tasked with advocating in favor of the claim, irrespective of its own opinion. The two agents then engage in a structured exchange, with the Persuader presenting arguments intended to influence the Persuadee’s stance. At the end of the conversation, the Persuadee provides a final decision and agreement score (see Turn $t$ in Figure 1). If at any point after the Persuader presents its first argument the Persuadee reports the maximum agreement score of Completely Support (5), the conversation is terminated early. The Persuadee is then prompted for its final decision without further arguments from the Persuader. This early stopping mechanism prevents unnecessary computation and ensures efficiency, as the model has no further room to be persuaded.
Throughout the interaction, system prompts guide response generation, with a dedicated decision prompt issued to the Persuadee in the final step. The overall approach is outlined in Algorithm 1, with further implementation details provided in Appendix B. The full set of prompts is available in Appendix E. To ensure a fair comparison of different Persuader models against the same Persuadee model, we fix the Persuadee’s initial message and agreement score for each claim $C$. This approach controls for variability in the Persuadee’s starting position, allowing us to isolate the influence of the Persuader’s arguments on opinion shifts.
@@ 3.2 Normalized Change in Agreement
To evaluate conversations in PMIYC, we introduce normalized change (NC) in agreement, a metric designed to quantify the persuasive impact of a conversation. In a $t$-turn dialogue, we compute the difference between the Persuadee’s initial agreement score, $sEE_{0}$, and its final agreement score, $sEE_{t}$. This difference represents the extent to which the Persuader has influenced the Persuadee’s stance. To ensure comparability across different initial agreement scores, we normalize the change using the following formulation:
$NC(c)=\begin{cases}\displaystyle\frac{sEE_{t}-sEE_{0}}{5-sEE_{0}},&\begin{array}[]{l}\text{if }sEE_{t}\geq sEE_{0},\\ \text{and }sEE_{0}\neq 5\end{array}\\[8.0pt] \displaystyle\frac{sEE_{t}-sEE_{0}}{sEE_{0}-1},&\text{ otherwise}\end{cases}$ (1)
The denominator ensures proper normalization by accounting for the range of possible agreement shifts. When the Persuadee’s agreement increases, the maximum possible increase is given by $(5-sEE_{0})$, while for a decrease, the maximum possible change is $(sEE_{0}-1)$. This normalization prevents bias toward Persuadees with initial scores closer to neutral or opposition, who would otherwise have greater potential for upward shifts. It also ensures that declines in agreement are properly accounted for, effectively penalizing the Persuader when persuasion efforts fail or lead to a stronger opposition stance.
Since the metric is normalized, its values always fall within the range $[-1,1]$. A value of 1 indicates a complete shift to agreement $(sEE_{t}=5)$, -1 signifies a full shift to opposition $(sEE_{t}=1)$, and 0 means no change. Negative values indicate cases where the persuasion attempt backfires, leading to stronger opposition rather than increased agreement. This metric provides a standardized way to compare conversations across different claims and agent pairings by measuring both effectiveness and susceptibility. Effectiveness refers to a model’s ability to persuade others and is computed as the average normalized change when the model acts as the Persuader. Susceptibility reflects how easily a model changes its stance and is measured as the average normalized change when the model serves as the Persuadee. Higher values indicate greater persuasive strength when the model is in the Persuader role and greater susceptibility to persuasion when it is in the Persuadee role.

### v3

# Multi-Turn Interactions in Large Language Models

@@ Abstract
Large Language Models (LLMs) demonstrate persuasive capabilities that rival human-level persuasion. While these capabilities can be used for social good, they also present risks of potential misuse. Beyond the concern of how LLMs persuade others, their own susceptibility to persuasion poses a critical alignment challenge, raising questions about robustness, safety, and adherence to ethical principles. To study these dynamics, we introduce Persuade Me If You Can (PMIYC), an automated framework for evaluating persuasiveness and susceptibility to persuasion in multi-agent interactions. Our framework offers a scalable alternative to the costly and time-intensive human annotation process typically used to study persuasion in LLMs. PMIYC automatically conducts multi-turn conversations between Persuader and Persuadee agents, measuring both the effectiveness of and susceptibility to persuasion. Our comprehensive evaluation spans a diverse set of LLMs and persuasion settings (e.g., subjective and misinformation scenarios). We validate the efficacy of our framework through human evaluations and demonstrate alignment with human assessments from prior studies. Through PMIYC, we find that Llama-3.3-70B and GPT-4o exhibit similar persuasive effectiveness, outperforming Claude 3 Haiku by 30%. However, GPT-4o demonstrates over 50% greater resistance to persuasion for misinformation compared to Llama-3.3-70B. These findings provide empirical insights into the persuasive dynamics of LLMs and contribute to the development of safer AI systems.  [^1] Code and data are available at https://beyzabozdag.github.io/PMIYC/.
@@ 1 Introduction
LLMs have demonstrated remarkable capabilities across various domains and are now integral to numerous real-world applications (OpenAI, 2024c; Anthropic, 2024). Among these capabilities, persuasion stands out as particularly noteworthy, with state-of-the-art LLMs exhibiting persuasive skills comparable to humans (Durmus et al., 2024; OpenAI, 2024c). Persuasive LLMs have been applied for social good, promoting public health and prosocial behaviors while adhering to ethical principles (Wang et al., 2019; Furumai et al., 2024; Karinshak et al., 2023; Kampik et al., 2018). However, as LLMs become increasingly persuasive, they also pose risks. These models can be exploited to not only manipulate individuals, spread misinformation, and influence public opinion in harmful ways (Singh et al., 2024; Salvi et al., 2024; Simchon et al., 2024; Hackenburg and Margetts, 2024), but also manipulate other agents to generate harmful outcomes, highlighting a growing need to understand their own susceptibility to persuasion. Such vulnerabilities are evident in jailbreaks, where LLMs are coerced into circumventing alignment principles, generating toxic content, or perpetuating harmful biases under adversarial inputs (Zeng et al., 2024; Xu et al., 2024). Persuasion susceptibility poses risks in many settings, such as customers coaxing models into bypassing financial or medical safeguards and rogue agents in multi-agent systems exploiting it to propagate misinformation or distort collective decision-making. This presents the challenge of persuasion susceptibility: models must resist harmful influence while remaining open to being convinced via justified rationale (Stengel-Eskin et al., 2024). To assess LLMs’ abilities to maintain this balance, we explore which conditions amplify the ability of and susceptibility to persuasion: (a) single-turn vs. multi-turn interactions, (b) subjective vs. misinformation claims, and (c) model family and size. As multi-agent systems become more prevalent, persuasion will become an essential skill in agent-to-agent interactions, and LLM persuasiveness and susceptibility to persuasion will become increasingly critical (Wu et al., 2023; Li et al., 2023; Hammond et al., 2025).
Figure 1: A $t$-turn interaction between a Persuader and a Persuadee in PMIYC. The Persuader addresses the Persuadee’s concerns, and the Persuadee reports its agreement scores.
@@
In this work, we introduce Persuade Me If You Can (PMIYC), an automated framework for evaluating persuasion in LLMs across diverse conversational setups. Building on recent studies (Durmus et al., 2024; Xu et al., 2024; Breum et al., 2023), our framework simulates multi-agent interactions where a Persuader attempts to convince a Persuadee to agree with a claim. PMIYC enables the Persuader and Persuadee to engage in multi-turn conversations, exchanging arguments, addressing counterpoints, and influencing each other’s views. Throughout these interactions, the Persuadee continuously assesses the Persuader’s arguments and self-reports its level of agreement, allowing us to track changes in opinion and measure both persuasive effectiveness (how successfully an LLM persuades) and susceptibility to persuasion (how easily an LLM changes its stance).
In PMIYC, each model takes turns acting as both the Persuader and the Persuadee against every other model, enables direct comparisons. Our evaluation spans a diverse set of models, including open-weight and closed-source models. We assess single-turn persuasion, where the Persuader has only one chance to convince the Persuadee, as well as multi-turn persuasion, which allows for back-and-forth exchanges and deeper argumentation. Additionally, we examine persuasion across different contexts, including subjective claims (eg. “Cultured/lab-grown meats should be allowed to be sold”) and misinformation claims (eg. “Canada is a part of the UK”), to evaluate how LLMs respond under varying conditions. However, PMIYC is flexible and can be extended to encompass other types of claims.
Our results reveal that persuasive effectiveness remains largely consistent across settings, with Llama-3.3-70B emerging as the most effective persuader, closely followed by GPT-4o. However, susceptibility to persuasion varies based on the context and the number of conversational turns. Notably, compared to other models, GPT-4o and GPT-4o-mini exhibit up to 50% stronger resistance to persuasion in misinformation contexts but show greater flexibility with subjective claims. Persuasive effectiveness is greater in the first two turns, and a Persuader becomes more effective when it more strongly agrees with the claim it advocates.
By modeling persuasion as an interactive and iterative process, PMIYC provides a more realistic setting for studying persuasion as it occurs in human-to-agent and agent-to-agent conversations. The reliability of our framework is validated through human evaluations and comparative analyses with existing studies. Our findings align closely with human judgments of persuasiveness, reinforcing the validity of PMIYC as a scalable alternative to human assessment while also allowing us to quantify levels of model susceptibility. Our key contributions are:
  • PMIYC, a framework for automatically evaluating both the persuasive effectiveness and susceptibility to persuasion of LLMs, especially for agent-to-agent communications.
@@
Through this research, we aim to deepen the understanding of persuasive capabilities and susceptibility of LLMs, and to emphasize the safety risks associated with persuasion and susceptibility in AI systems. By providing a scalable and automated evaluation framework, we aim to contribute to the development of safer, more transparent, and ethically responsible LLMs for real-world applications.
@@ 2 Related Work
Persuasive Language Models. Recent studies suggest that LLMs exhibit human-like persuasiveness (Durmus et al., 2024; OpenAI, 2024c; Huang and Wang, 2023). Durmus et al. (2024)’s and OpenAI (2024c)’s experiments show that their models are as persuasive as humans. While some studies leverage the persuasive abilities of LLMs for social good (Karinshak et al., 2023; Costello et al., 2024), others warn of risks like microtargeting, political influence, and ethical concerns (Salvi et al., 2024; Simchon et al., 2024; Hackenburg and Margetts, 2024). These findings underscore the growing urgency of understanding and regulating persuasive AI systems.
@@
Evaluating Persuasion. Prior research has explored various methods to measure the persuasiveness of LLMs. Durmus et al. (2024) uses human evaluators to rank their agreement with a subjective claim before and after reading the model-generated arguments. While this approach provides valuable insights into human preferences, it is not highly scalable. Similarly, OpenAI (2024c) uses human annotators by having them rank arguments and perform pairwise comparisons. They also explore automated evaluation through their MakeMeSay and MakeMePay role-playing frameworks. These frameworks, while innovative, are highly context-dependent and may not generalize well to real-world persuasion scenarios. Singh et al. (2024) introduce PersuasionArena and the task of transsuasion, where LLMs rewrite low-engagement tweets into more persuasive versions. Although it is an automated framework for measuring persuasion, its focus is less conversational and does not fully address the interactive and manipulative aspects of persuasion. Breum et al. (2023) develop a Convincer-Skeptic scenario, in which an LLM attempts to persuade another LLM in the domain of climate change. Zhou et al. (2024) propose SOTOPIA, a role-play-based platform for evaluating social intelligence, which includes persuasive elements but does not directly quantify persuasion effectiveness or resistance and relies on LLM-as-a-judge scoring. Other studies have attempted to score persuasive language in pairs of LLM-generated text using a regression model (Pauli et al., 2024). Yet, these efforts focus on short text excerpts, limiting their applicability to longer, multi-turn persuasion.
@@
LLM Susceptibility to Persuasion. While LLMs generating persuasive content raises ethical concerns, their susceptibility to persuasion is equally important. Some jailbreaking studies show that LLMs can be persuaded to generate harmful content (Zeng et al., 2024) or accept and reinforce misinformation (Xu et al., 2024). Furthermore, recent studies have shown that multi-turn jailbreak attempts can be particularly effective in circumventing safeguards (Xu et al., 2024; Li et al., 2024; Russinovich et al., 2024). Inspired by these findings, we extend our framework to the misinformation domain and employ a dual evaluation approach that captures not only persuasive effectiveness but also a model’s susceptibility to persuasion.
@@
Overall, prior research has primarily focused on single-turn persuasive attacks on LLMs, human-in-the-loop evaluation methods, and less conversational or highly constrained scenarios, with limited exploration of multi-turn agent-to-agent persuasion. In contrast, we explore how LLMs interact with each other in a highly conversational setting across different domains and multi-turn scenarios, encompassing both beneficial and harmful persuasion.
@@ 3 Methodology
PMIYC is our framework for evaluating both the persuasive effectiveness of LLMs and their susceptibility to persuasion. [^2] Code adapted from NegotiationArena Bianchi et al. (2024). Experiments consist of multi-agent interactions, where one agent acts as a Persuader (ER) and the other acts as a Persuadee (EE). The two engage in a structured exchange, where the Persuader agent is given the objective of convincing the Persuadee to agree with a given claim. The Persuadee engages with the Persuader and reassesses its stance on the claim throughout the conversation. While using LLM-judges to assess persuasiveness may appear more straightforward and similar in nature, we demonstrate in Appendix E that this approach is unreliable, further motivating the need for PMIYC.
@@ 3.1 Persuasive Conversation Generation
Input: Number of turns $t$
Output: Full conversation history CHCH, turn-level agreement score history of the persuadee sEEsEE
0.5em $CH\leftarrow[]$, $sEE\leftarrow[]$
0.5em for $turn_{i}\in\{1,2,\ldots,t-1\}$ do
  if $turn_{i}$ is odd then
     // $E$ turn
     $msgEE_{i},sEE_{i}=EE(CH,sEE)$;
     Append $msgEE_{i}$ to CHCH;
     Append $sEE_{i}$ to sEEsEE;
     if $sEE_{i}=5$ and $turn_{i}\neq 1$ then
        // $E$ fully persuaded
        break;
 else
     // ER turn
     $msgER_{i}=ER(CH)$;
     Append $msgER_{i}$ to CHCH;
// $E$’s final decision (Turn t)
$msg_{t},sEE_{t}=\text{FinalDecision}(CH,sEE)$;
Append $msg_{t}$ to CHCH;
Append $sEE_{t}$ to sEEsEE;
Algorithm 1 Persuasive Conversation Generation Algorithm
Conversation generation involves simulating a dialogue between two agents: a Persuader and a Persuadee. The objective is to track how the Persuadee’s stance on a given claim $C$ evolves over the course of a $t$-turn interaction. The conversation begins with the Persuadee expressing its initial agreement with the claim by providing both a message and an agreement score (see Turn 1 in Figure 1). The agreement score quantifies the model’s stance on claim $C$ using a five-point Likert scale: Completely Oppose (1), Oppose (2), Neutral (3), Support (4), and Completely Support (5). This scale enables a more granular analysis of opinion shifts beyond a binary agree/disagree classification. The use of a Likert scale to measure shifts in mental states or beliefs is motivated by Durmus et al. (2024) and supported by prior work in psychology and social sciences, where similar scales are commonly employed to assess persuasiveness and belief change in human subjects (Thomas et al., 2019; Kodapanakkal et al., 2022).
After observing the Persuadee’s initial stance, the Persuader is tasked with advocating in favor of the claim, irrespective of its own opinion. The two agents then engage in a structured exchange, with the Persuader presenting arguments intended to influence the Persuadee’s stance. At the end of the conversation, the Persuadee provides a final decision and agreement score (see Turn $t$ in Figure 1). If at any point after the Persuader presents its first argument, the Persuadee reports the maximum agreement score of Completely Support (5), the conversation is terminated early. The Persuadee is then prompted for its final decision without further arguments from the Persuader. This early stopping mechanism prevents unnecessary computation and ensures efficiency, as the model has no further room to be persuaded.
Throughout the interaction, system prompts guide response generation, with a dedicated decision prompt issued to the Persuadee in the final step. The overall approach is outlined in Algorithm 1, with further implementation details provided in Appendix B. The full set of prompts is available in Appendix K. When pairing different Persuader models with a Persuadee, we standardize the Persuadee’s initial message and agreement score for each claim $C$ for all Persuadees. This ensures fair comparisons between different Persuaders by reducing the variability of the Persuadee’s starting positions.
@@ 3.2 Normalized Change in Agreement
To evaluate conversations in PMIYC, we introduce normalized change in agreement (NCA), a metric designed to quantify the persuasive impact of a conversation. In a $t$-turn dialogue, we compute the difference between the Persuadee’s initial agreement score, $sEE_{0}$, and its final agreement score, $sEE_{t}$. This difference represents the extent to which the Persuader has influenced the Persuadee’s stance. To ensure comparability across different initial agreement scores, we normalize the change using the following formulation:
$NCA(c)=\begin{cases}\displaystyle\frac{sEE_{t}-sEE_{0}}{5-sEE_{0}},&\begin{array}[]{l}\text{if }sEE_{t}\geq sEE_{0},\\ \text{and }sEE_{0}\neq 5\end{array}\\[8.0pt] \displaystyle\frac{sEE_{t}-sEE_{0}}{sEE_{0}-1},&\text{ otherwise}\end{cases}$ (1)
The denominator ensures proper normalization by accounting for the range of possible agreement shifts. When the Persuadee’s agreement increases, the maximum possible increase is given by $(5-sEE_{0})$, while for a decrease, the maximum possible change is $(sEE_{0}-1)$. This normalization prevents bias toward Persuadees with initial scores closer to neutral or opposition, who would otherwise have more potential for a positive score change. It also ensures that declines in agreement are properly accounted for, effectively penalizing the Persuader when persuasion efforts fail or lead to a stronger opposition stance.
Since the metric is normalized, its values always fall within the range [1,1][-1,1]. A value of 1 indicates a complete shift to agreement $(sEE_{t}=5)$, -1 signifies a full shift to opposition $(sEE_{t}=1)$, and 0 means no change. Negative values indicate cases where the persuasion attempt backfires, leading to stronger opposition rather than increased agreement. This metric provides a standardized way to compare conversations across different claims and agent pairings by measuring both effectiveness and susceptibility. Effectiveness refers to a model’s ability to persuade others and is computed as the average normalized change when the model acts as the Persuader. Susceptibility reflects how easily a model changes its stance and is measured as the average normalized change when the model serves as the Persuadee. Higher values indicate greater persuasive strength when the model is in the Persuader role and greater susceptibility to persuasion when it is in the Persuadee role.
Our use of agreement score differences as a proxy for belief change aligns with established definitions of persuasion in the social sciences, which characterize it as the act of shaping, reinforcing, or altering an individual’s mental state or response (Miller, 1980; O’Keefe, 2002). Crucially, this framing captures both shifts from opposition to support and strengthening of existing supportive attitudes, underscoring that persuasion is not limited to changing minds but also includes reinforcing existing beliefs. By quantifying this change regardless of the starting point, our metric provides a robust and theoretically grounded measure of persuasive impact. More details on the use of normalized scores, along with comparative absolute results, are provided in Appendix I. While a multi-faceted evaluation pipeline is ideal for capturing belief change, we find that self-reported agreement serves as a reliable and scalable proxy within PMIYC. Action-based MCQ tasks (§ 5.3.2), LLM-as-judge assessments (§ D.1), and human annotations (§ 5.3.1) consistently align with self-reports.